NvidiaIntely (nvidiaintely):

Determine the molecular formula for glucose, with a molar mass of 180 grams, if the empirical formula for glucose is CH2O. CH2O C2H4O2 C3H6O3 C6H12O6

OpenStudy (welshfella):

take the empirical formula CH2O adding the atomic masses 12 + 2*1 + 16 = 30 Now divide 180 by 30

OpenStudy (welshfella):

what do you get?

OpenStudy (welshfella):

that gives you the factor you'll use to find the molecular formula

OpenStudy (welshfella):

180 / 30 = ?

NvidiaIntely (nvidiaintely):

Let me see.. P.S. sorry I was gn for a moment.

OpenStudy (jiteshmeghwal9):

Molecular formula=\((empirical formula)_n\)

NvidiaIntely (nvidiaintely):

C = 12 / MM CH2O (30) = 40 % H = 2/30 = 6.67% ; O = 16 / 30 = 53.33 % C = 40% of 180g = 76g or 6 moles of C H = 6.67 % of 180 = 12g or 12 moles of H O = 53.93% of 180 = 98 g or 6 moles of O

OpenStudy (jiteshmeghwal9):

n=(molar mass)/(empirical mass)

OpenStudy (jiteshmeghwal9):

Empirical mass=12*1+1*2+1*16=30 Molar mass=180 n=6

OpenStudy (jiteshmeghwal9):

Molecular formula=\((CH_2O)_n\)

OpenStudy (jiteshmeghwal9):

Molecular formula=\((CH_2O)_6\)

OpenStudy (welshfella):

yes n = 180/30 = 6

OpenStudy (kevin):

You need to find its mole ratio mol = mass/ Mr 180/(12+2+16)=180/30 = 6 Now you need to multiply it with the empiric formula :)

NvidiaIntely (nvidiaintely):

So @welshfella, the answer is C6H12O6.

OpenStudy (welshfella):

In this case there is no need to work out the percentages of each of the elements

OpenStudy (welshfella):

yes

NvidiaIntely (nvidiaintely):

YEAH!!!! I DID IT! Thanks guys!

OpenStudy (welshfella):

yw

OpenStudy (kevin):

urwel